# Parameter-Efficient Fine-Tuning for Continual Learning: A Neural Tangent Kernel Perspective

### v2

@@ Appendix H Pre-trained Weight Matters
TABLE XV: The evolution of incremental top-1 accuracy during the full fine-tuning process between the original ViT-B/16 model and an enhanced version incorporating three auxiliary subnetworks, highlighting the changes throughout training.
Network Incremental Top-1 Accuracy
@@ Appendix I More Visualizations
[…]
@@
Dataset $\mathcal{C}^{base}$ $\mathcal{N}_{base}$ $\mathcal{C}^{inc}$Inc. Shots Resolution
@@
TABLE XVII: The evolution of incremental top-1 accuracy for different datasets under the FSCIL setting, using pre-trained weights from ImageNet-21K. Bold segments indicate optimal results, while underlined segments denote suboptimal results.
FSCIL Methods Incremental Top-1 Accuracy
@@
CEC [104] CIFAR100 80.53 76.43 73.47 69.88 67.58 65.28 63.87 61.93 59.86 - -
ALICE [62] CIFAR100 82.42 73.42 70.74 67.44 65.87 63.68 62.32 60.56 58.59 - -
NTK-CL (Ours) CIFAR100 93.92 91.42 90.61 89.21 89.08 88.39 88.38 87.84 86.42 - -
CEC [104] miniImageNet 94.82 92.34 89.69 87.84 86.83 84.53 82.28 81.51 81.05 - -
ALICE [62] miniImageNet 92.72 90.83 88.41 86.89 85.70 83.46 81.66 80.60 80.09 - -
NTK-CL (Ours) miniImageNet 97.67 97.20 95.49 95.09 95.00 94.17 92.99 92.84 92.75 - -
CEC [104] CUB200 84.51 82.68 80.47 76.57 76.47 74.77 74.76 74.08 72.72 72.37 71.55
ALICE [62] CUB200 77.65 69.71 68.66 68.48 67.92 66.44 65.91 64.68 64.60 64.35 63.83
NTK-CL (Ours) CUB200 89.87 88.22 87.60 86.21 85.07 84.80 84.56 84.51 84.45 84.30 84.28
TABLE XVIII: The evolution of incremental top-1 accuracy for different datasets under the Imbalanced CIL setting, utilizing the pre-trained weight derived from the ImageNet-21K. The suffix ’-LFS’ denotes uniform partitioning of all classes into $N$ tasks for incremental training from scratch, while the suffix ’-LFH’ involves initial training on the first half of classes followed by incremental learning of the remaining classes divided into $N$ tasks. Bold segments indicate optimal results, while underlined segments denote suboptimal results.
Imbalanced CIL Methods Incremental Top-1 Accuracy
@@
LT-CIL-LFS [54] CIFAR100-LT 83.10 72.25 69.87 65.65 63.66 59.93 59.14 57.54 56.24 55.45 -
GR-LFS [27] CIFAR100-LT 84.60 75.40 70.20 65.12 62.46 58.88 58.26 56.11 55.37 54.72 -
NTK-CL-LFS (Ours) CIFAR100-LT 82.60 78.70 77.27 74.25 72.76 71.26 71.12 69.97 69.77 69.40 -
LT-CIL-LFS [54] ImageNetSubset-LT 94.40 91.00 89.80 89.55 89.31 87.46 86.88 84.15 83.59 83.33 -
GR-LFS [27] ImageNetSubset-LT 96.00 93.90 92.47 92.20 92.08 90.49 90.36 86.67 86.44 86.07 -
NTK-CL-LFS (Ours) ImageNetSubset-LT 96.40 93.90 92.80 92.80 92.59 90.97 90.46 88.53 88.48 88.22 -
LT-CIL-LFH [54] CIFAR100-LT 62.64 57.38 51.98 54.79 56.75 55.06 55.31 54.49 54.14 54.15 54.15
GR-LFH [27] CIFAR100-LT 65.06 62.29 58.90 59.90 61.11 59.77 58.50 58.85 57.81 57.39 56.54
NTK-CL-LFH (Ours) CIFAR100-LT 83.18 77.90 76.54 76.54 78.38 77.24 76.31 76.82 76.88 76.62 76.43
LT-CIL-LFH [54] ImageNetSubset-LT 90.55 90.08 87.95 86.98 87.97 86.70 82.43 83.41 84.28 82.99 82.44
GR-LFH [27] ImageNetSubset-LT 92.95 90.72 90.88 92.19 91.93 91.15 87.08 87.27 87.06 87.08 86.68
NTK-CL-LFH (Ours) ImageNetSubset-LT 94.04 93.44 93.81 94.45 94.35 94.11 90.63 90.55 90.59 89.98 90.25

### v3

@@ Appendix H Pre-trained Weight Matters
TABLE XV: Evolution of incremental top-1 accuracy (%) on CIFAR100 during full fine-tuning, comparing the original ViT-B/16 model with an enhanced variant incorporating three auxiliary subnetworks.
Network Incremental Top-1 Accuracy (%)
@@ Appendix I More Visualizations
Code Snippet
1. Inputs: $\mathcal{L}_{dis},\mathcal{L}_{orth},\mathcal{L}_{reg};\ \eta\in[\eta_{\min},\eta_{\max}]([0.1,0.5]),\upsilon\in[\upsilon_{\min},\upsilon_{\max}]([1e-5,1e-3]),\lambda\in[\lambda_{\min},\lambda_{\max}]([1e-5,1e-3]);\ \beta=0.95$.
2. Initialization: $\mu_{dis},\mu_{orth},\mu_{reg}\leftarrow 0;\ \nu_{dis},\nu_{orth},\nu_{reg}\leftarrow 0;\ \eta\leftarrow\eta_{\min},\ \upsilon\leftarrow\upsilon_{\min},\ \lambda\leftarrow\lambda_{\min}$.
3. For each training iteration $t$ do:
Update moving averages: $\mu_{dis}\leftarrow\beta\mu_{dis}+(1-\beta)l_{dis}^{(t)};\ \nu_{dis}\leftarrow\beta\nu_{dis}+(1-\beta)\left(l_{dis}^{(t)}\right)^{2}$; (repeat for $\mu_{orth},\nu_{orth},\mu_{reg},\nu_{reg}$).
Compute standard deviations: $\sigma_{dis}\leftarrow\sqrt{\max(\nu_{dis}-\mu_{dis}^{2},\ 0)}$; (repeat for $\sigma_{orth},\sigma_{reg}$).
Normalize deviations: $\delta_{dis}\leftarrow\dfrac{l_{dis}^{(t)}-\mu_{dis}}{\sigma_{dis}}$; (repeat for $\delta_{orth},\delta_{reg}$).
Non-linear squashing: $\epsilon_{dis}\leftarrow\tanh(\delta_{dis})$; (repeat for $\epsilon_{orth},\epsilon_{reg}$).
Compute target weights: $\eta^{target}\leftarrow\epsilon_{dis}\cdot(\eta_{\max}-\eta_{\min})$; (repeat for $\upsilon^{target},\lambda^{target}$).
EMA smoothing of weights: $\eta\leftarrow\beta\eta+(1-\beta)\eta^{target}$; (repeat for $\upsilon,\lambda$).
Clip to valid range: $\eta\leftarrow\mathrm{clip}(\eta,\eta_{\min},\eta_{\max})$; (repeat for $\upsilon,\lambda$).
4. Return: $\eta,\upsilon,\lambda$.
@@ 3. For each training iteration $t$ do:
Dataset $\mathcal{C}^{base}$ $\mathcal{N}_{base}$ $\mathcal{C}^{inc}$ #Inc. Shots Resolution
@@
TABLE XVII: Evolution of incremental top-1 accuracy (%) for different datasets under the FSCIL setting, using ImageNet-21K pre-trained weights. Bold segments indicate the best results, while underlined segments denote suboptimal outcomes.
FSCIL Methods Incremental Top-1 Accuracy (%)
@@
CEC [104] CIFAR100 80.71 $\pm$ 0.51 76.34 $\pm$ 0.27 73.46 $\pm$ 0.19 70.03 $\pm$ 0.40 67.39 $\pm$ 0.31 65.35 $\pm$ 0.23 64.04 $\pm$ 0.20 62.12 $\pm$ 0.39 59.69 $\pm$ 0.36 - -
ALICE [62] CIFAR100 82.23 $\pm$ 0.35 73.34 $\pm$ 0.31 70.81 $\pm$ 0.35 67.33 $\pm$ 0.26 65.92 $\pm$ 0.26 63.79 $\pm$ 0.21 62.37 $\pm$ 0.28 60.61 $\pm$ 0.28 58.60 $\pm$ 0.24 - -
NTK-CL (Ours) CIFAR100 94.08 $\pm$ 0.47 91.48 $\pm$ 0.12 90.54 $\pm$ 0.15 89.16 $\pm$ 0.25 89.14 $\pm$ 0.29 88.34 $\pm$ 0.23 88.22 $\pm$ 0.28 87.88 $\pm$ 0.16 86.30 $\pm$ 0.32 - -
CEC [104] miniImageNet 94.92 $\pm$ 0.25 92.36 $\pm$ 0.17 89.52 $\pm$ 0.29 87.98 $\pm$ 0.17 86.75 $\pm$ 0.32 84.62 $\pm$ 0.15 82.23 $\pm$ 0.26 81.50 $\pm$ 0.10 80.91 $\pm$ 0.16 - -
ALICE [62] miniImageNet 92.81 $\pm$ 0.37 90.65 $\pm$ 0.42 88.57 $\pm$ 0.20 86.88 $\pm$ 0.31 85.80 $\pm$ 0.12 83.43 $\pm$ 0.22 81.81 $\pm$ 0.30 80.77 $\pm$ 0.39 80.09 $\pm$ 0.25 - -
NTK-CL (Ours) miniImageNet 97.75 $\pm$ 0.15 97.04 $\pm$ 0.32 95.30 $\pm$ 0.25 95.01 $\pm$ 0.26 94.92 $\pm$ 0.20 94.29 $\pm$ 0.18 93.09 $\pm$ 0.10 92.84 $\pm$ 0.21 92.81 $\pm$ 0.30 - -
CEC [104] CUB200 84.67 $\pm$ 0.30 82.68 $\pm$ 0.20 80.52 $\pm$ 0.27 76.55 $\pm$ 0.25 76.47 $\pm$ 0.19 74.74 $\pm$ 0.11 74.69 $\pm$ 0.25 74.05 $\pm$ 0.25 72.58 $\pm$ 0.23 72.26 $\pm$ 0.26 71.56 $\pm$ 0.15
ALICE [62] CUB200 77.47 $\pm$ 0.25 69.73 $\pm$ 0.26 68.70 $\pm$ 0.22 68.64 $\pm$ 0.21 67.89 $\pm$ 0.12 66.25 $\pm$ 0.38 66.00 $\pm$ 0.17 64.77 $\pm$ 0.16 64.62 $\pm$ 0.17 64.22 $\pm$ 0.24 63.73 $\pm$ 0.32
NTK-CL (Ours) CUB200 89.81 $\pm$ 0.26 88.15 $\pm$ 0.29 87.67 $\pm$ 0.24 86.24 $\pm$ 0.13 84.94 $\pm$ 0.32 84.78 $\pm$ 0.25 84.57 $\pm$ 0.34 84.49 $\pm$ 0.24 84.40 $\pm$ 0.19 84.38 $\pm$ 0.20 84.22 $\pm$ 0.22
TABLE XVIII: Evolution of incremental top-1 accuracy (%) for different datasets under the Imbalanced CIL setting, utilizing the pre-trained weight derived from the ImageNet-21K. The suffix ’-LFS’ denotes uniform partitioning of all classes into NN tasks for incremental training from scratch, while the suffix ’-LFH’ involves initial training on the first half of classes followed by incremental learning of the remaining classes divided into NN tasks. Bold segments indicate optimal results, while underlined segments denote suboptimal results.
Imbalanced CIL Methods Incremental Top-1 Accuracy (%)
@@
LT-CIL-LFS [54] CIFAR100-LT 83.21 $\pm$ 0.32 72.42 $\pm$ 0.35 69.86 $\pm$ 0.18 65.57 $\pm$ 0.27 63.86 $\pm$ 0.33 59.91 $\pm$ 0.09 59.15 $\pm$ 0.18 57.40 $\pm$ 0.30 56.31 $\pm$ 0.13 55.30 $\pm$ 0.40 -
GR-LFS [27] CIFAR100-LT 84.59 $\pm$ 0.18 75.56 $\pm$ 0.20 70.15 $\pm$ 0.21 65.29 $\pm$ 0.30 62.32 $\pm$ 0.31 58.82 $\pm$ 0.21 58.16 $\pm$ 0.28 56.15 $\pm$ 0.18 55.38 $\pm$ 0.22 54.62 $\pm$ 0.18 -
NTK-CL-LFS (Ours) CIFAR100-LT 82.67 $\pm$ 0.23 78.71 $\pm$ 0.14 77.24 $\pm$ 0.22 74.05 $\pm$ 0.40 72.63 $\pm$ 0.21 71.11 $\pm$ 0.33 71.07 $\pm$ 0.20 69.95 $\pm$ 0.20 69.90 $\pm$ 0.20 69.57 $\pm$ 0.23 -
LT-CIL-LFS [54] ImageNetSubset-LT 94.29 $\pm$ 0.12 90.85 $\pm$ 0.27 89.75 $\pm$ 0.26 89.47 $\pm$ 0.28 89.25 $\pm$ 0.16 87.45 $\pm$ 0.16 87.01 $\pm$ 0.34 84.22 $\pm$ 0.17 83.49 $\pm$ 0.38 83.39 $\pm$ 0.21 -
GR-LFS [27] ImageNetSubset-LT 96.04 $\pm$ 0.29 93.83 $\pm$ 0.15 92.52 $\pm$ 0.29 92.24 $\pm$ 0.21 92.16 $\pm$ 0.20 90.38 $\pm$ 0.23 90.35 $\pm$ 0.20 86.67 $\pm$ 0.20 86.45 $\pm$ 0.22 86.24 $\pm$ 0.29 -
NTK-CL-LFS (Ours) ImageNetSubset-LT 96.30 $\pm$ 0.31 93.90 $\pm$ 0.18 92.94 $\pm$ 0.18 92.87 $\pm$ 0.23 92.60 $\pm$ 0.10 91.00 $\pm$ 0.15 90.51 $\pm$ 0.23 88.57 $\pm$ 0.18 88.46 $\pm$ 0.22 88.06 $\pm$ 0.38 -
LT-CIL-LFH [54] CIFAR100-LT 62.74 $\pm$ 0.14 57.40 $\pm$ 0.14 52.08 $\pm$ 0.26 54.87 $\pm$ 0.20 56.58 $\pm$ 0.26 55.22 $\pm$ 0.21 55.37 $\pm$ 0.15 54.30 $\pm$ 0.30 54.21 $\pm$ 0.23 54.15 $\pm$ 0.10 54.16 $\pm$ 0.12
GR-LFH [27] CIFAR100-LT 65.06 $\pm$ 0.14 62.41 $\pm$ 0.22 59.08 $\pm$ 0.28 59.76 $\pm$ 0.27 61.09 $\pm$ 0.11 59.75 $\pm$ 0.18 58.52 $\pm$ 0.12 59.04 $\pm$ 0.31 57.81 $\pm$ 0.14 57.30 $\pm$ 0.25 56.66 $\pm$ 0.20
NTK-CL-LFH (Ours) CIFAR100-LT 83.13 $\pm$ 0.12 77.97 $\pm$ 0.12 76.55 $\pm$ 0.10 76.35 $\pm$ 0.33 78.32 $\pm$ 0.18 77.42 $\pm$ 0.33 76.43 $\pm$ 0.22 76.89 $\pm$ 0.09 76.87 $\pm$ 0.14 76.78 $\pm$ 0.22 76.42 $\pm$ 0.10
LT-CIL-LFH [54] ImageNetSubset-LT 90.68 $\pm$ 0.24 90.25 $\pm$ 0.33 87.97 $\pm$ 0.14 86.97 $\pm$ 0.16 87.87 $\pm$ 0.20 86.78 $\pm$ 0.18 82.56 $\pm$ 0.25 83.54 $\pm$ 0.20 84.24 $\pm$ 0.17 83.17 $\pm$ 0.28 82.48 $\pm$ 0.19
GR-LFH [27] ImageNetSubset-LT 93.08 $\pm$ 0.31 90.55 $\pm$ 0.30 90.79 $\pm$ 0.19 92.12 $\pm$ 0.12 91.99 $\pm$ 0.16 91.09 $\pm$ 0.10 86.94 $\pm$ 0.29 87.24 $\pm$ 0.16 86.96 $\pm$ 0.14 86.93 $\pm$ 0.19 86.77 $\pm$ 0.22
NTK-CL-LFH (Ours) ImageNetSubset-LT 94.16 $\pm$ 0.24 93.37 $\pm$ 0.14 93.72 $\pm$ 0.20 94.37 $\pm$ 0.16 94.19 $\pm$ 0.19 94.30 $\pm$ 0.31 90.73 $\pm$ 0.19 90.58 $\pm$ 0.11 90.75 $\pm$ 0.24 90.04 $\pm$ 0.16 90.18 $\pm$ 0.14